# LeanQuant: Accurate and Scalable Large Language Model Quantization with Loss-error-aware Grid

### v2

@@ 5 Related Works
Iterative Loss-error-based Compression Optimal Brain Damage (LeCun et al., 1989) introduced a saliency-score-based iterative pruning algorithm for neural networks, and Optimal Brain Surgeon (Hassibi & Stork, 1992; Hassibi et al., 1993) extended it to apply a weight update to compensate for the error introduced in each iteration. These methods inspired a number of works on model pruning (Guo et al., 2016; Singh & Alistarh, 2020; Yu et al., 2022) and weight quantization (Li et al., 2021; Frantar & Alistarh, 2022; Frantar et al., 2022).
Efficient LLM Inference LLM inference is computationally and memory demanding, and existing works accelerate inference and reduce memory requirements through post-training weight quantization (Dettmers et al., 2022; Lin et al., 2024; Frantar et al., 2022; Chee et al., 2024; Kim et al., 2023; Shao et al., 2024; Egiazarian et al., 2024; Tseng et al., 2024), pruning (Frantar & Alistarh, 2023; Ashkboos et al., 2024), weight-activation quantization (Xiao et al., 2023), offloading Sheng et al. (2023), etc.
@@ 6 Conclusion
Appendix A LeanQuant-Exact
@@
3:$\hat{\mathbf{w}}\leftarrow\mathbf{0}_{c}$
4:$\mathbf{H}^{-1}\leftarrow(2\mathbf{X}\mathbf{X}^{\top})^{-1}$
@@
6:   $\mathcal{G}\leftarrow\operatorname*{arg\,min}\limits_{\mathcal{G}:\left|\mathcal{G}\right|=2^{b}}\big{(}\mathrm{diag}(\mathbf{H}^{-1})^{-p}\big{)}^{\top}\big{|}\mathrm{quant}_{\textit{nu}}(\mathbf{w},\mathcal{G})-\mathbf{w}\big{|}^{2}$triangleright E. 6
@@
8:   $S,Z\leftarrow\operatorname*{arg\,min}\limits_{(S,Z)\in\mathbb{S}}\big{(}\mathrm{diag}(\mathbf{H}^{-1})^{-p}\big{)}^{\top}\big{|}\mathrm{quant}_{\textit{aff}}(\mathbf{w},\mathcal{S},Z)-\mathbf{w}\big{|}^{2}$triangleright E. 7
@@
18:   $\mathbf{w}\leftarrow\mathbf{w}-\frac{\mathbf{H}^{-1}_{:,i}}{\mathbf{H}^{-1}_{i,i}}\big{(}w_{i}-\hat{w}_{i}\big{)}$
19:   $\mathbf{H}^{-1}\leftarrow\mathbf{H}^{-1}-\frac{\mathbf{H}^{-1}_{:,i}\mathbf{H}^{-1}_{i,:}}{\mathbf{H}^{-1}_{i,i}}$
@@
A.1 BERT Experiments with LeanQuant-Exact
We compare the performance of BERT models (Devlin et al., 2018), quantized with OBQ (Frantar & Alistarh, 2022) and LeanQuantnu-Exact, on the SQuAD dataset (Rajpurkar et al., 2016). We quantize the 12-layer BERT-base (Devlin et al., 2018) and the 3-layer BERT-3 variant from Kurtic et al. (2022) to 3 and 4 bits. OBQ and LeanQuant-Exact are calibrated using 1024 samples from the training set, and the F1 score is reported on the test set.
@@ A.1 BERT Experiments with LeanQuant-Exact
Appendix B Discussion on Error Accumulation During Iterative Quantization
LeanQuant prevents drastic increase to the task loss by learning the quantization grid for better preservation of the precision of outlier inverse diagonals. However, since the not-yet-quantized weights will shift during the iterative quantization process and the quantization grid is fixed beforehand, one potential problem arises: the quantization grid is no longer well-aligned with the outliers after certain iterations. Fortunately, this is not a problem in practice. The loss-error-awareness property of LeanQuant grids prevents high-norm weight perturbations $\bm{\delta}_{i}$ (Equation 3) from ocurring, hence the weights do not shift by much during the iterations. Furthermore, no new inverse-diagonal outliers will arise during the iterative quantization process, since the remaining inverse diagonals only decrease in magnitude towards zero after each column and row removal (Equation 4).
Appendix C Experiment Details
Baseline Reproduction
Perplexity Evaluations
We follow the perplexity evaluation procedure described by (Frantar et al., 2022): sequences from the test set of the WikiText2 and C4 datasets (Merity et al., 2016; Raffel et al., 2020) are concatenated into 128 sequences of length 2048 tokens for perplexity testing.
Accuracy Evaluations
We use lm-evaluation-harness (Gao et al., 2023) for evaluating zero-shot accuracy on tasks. The task names we evaluate are lambada, ai2_arc, winogrande, piqa, hellaswag, mmlu.
Appendix D Perplexity Evaluations
The perplexity evaluation results on WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020) for quantized models are presented in Table 7.
Appendix E LLM-as-a-Judge
Appendix F Accuracy Results on More Models
The zero-shot accuracy results on benchmarks for quantized LLaMA-7B, LLaMA-13B, Llama-2-7B (Touvron et al., 2023a; b) are presented in Table 8.
Table 7: Perplexity evaluations of Llama models under different quantization methods and bit widths. The results of GPTQ, AWQ, OmniQuant are from Shao et al. (2024), and the results of SqueezeLLM are from Kim et al. (2023). The official SqueezeLLM code does not support 2-bit quantization, and we report the available results from Kim et al. (2023).
@@ Appendix F Accuracy Results on More Models
Non-uniform SqueezeLLM 4.05 76.56 46.25 69.53 75.22 32.98 37.19 42.18 44.00 59.29 78.62 71.82 57.60
LeanQuantnu 4.05 76.39 45.05 71.55 76.48 34.76 38.77 46.12 47.18 59.29 78.78 73.24 58.87
@@
Appendix G Quantization Time Cost
The time cost of LeanQuant for different models and configurations are presented in Table 9.
@@ Appendix G Quantization Time Cost
Appendix H Ablation Study
@@
Appendix I Loss Error Comparison

### v3

@@ 5 Related Works
Iterative Loss-error-based Compression Optimal Brain Damage (obd) introduced a saliency-score-based iterative pruning algorithm for neural networks, and Optimal Brain Surgeon (second_obs; obs) extended it to apply a weight update to compensate for the error introduced in each iteration. These methods inspired a number of works on model pruning (dynamic_surgery; woodfisher; cbs) and weight quantization (brecq; obq; gptq).
Efficient LLM Inference LLM inference is computationally and memory demanding, and existing works explore improving inference efficiency through quantization (llmint8; awq; gptq; quip; sqllm; omniquant; aqlm; quipsharp), pruning (sparsegpt; slicegpt), weight-activation quantization (smoothquant), offloading flexgen, etc. A survey of more relevant literature can be found in Appendix M.
@@ 6 Conclusion
Acknowledgements
This work was supported by National Science Foundation SHF-2211815 and Ken Kennedy Institute Cluster Grants.
@@ Acknowledgements
Appendix A Explanations on Quantization Grid
Figure 3: Comparison of affine (left) and non-uniform (right) 2-bit quantization grids applied to the weights in the first MLP-down layer of Llama-3-8B. The affine grid uses evenly spaced quantization grid points between the minimum and maximum weights. In contrast, the non-uniform grid allows grid points to be placed flexibly, as their positions are stored in a look-up table. This enables finer quantization in dense regions and coarser quantization in sparse regions, better aligning with the weight distribution and reducing quantization error.
In the context of quantization, a grid is a predefined set of values representing the possible quantized outputs for full-precision parameters. During quantization, each full-precision parameter is mapped to its nearest grid point on the quantization grid. For example, in a 2-bit quantization scheme with grid points $\{-1.0,-0.33,0.33,1.0\}$, a floating-point weight of 0.25 would be assigned to 0.33, the closest grid point.
Affine Quantization Grid
An affine quantization grid distributes points uniformly across the range of the weights being quantized. The dynamic range of the weights, defined as $W_{\mathrm{min}},W_{\mathrm{max}}]$, determines the spacing of the grid points. For example, if $W_{\mathrm{min}},W_{\mathrm{max}}]=[-1.0,1.0]$ in a 2-bit quantization setting, the grid points would be evenly spaced at 1.0,0.33,0.33,1.0-1.0,-0.33,0.33,1.0. This uniform distribution is computationally simple and widely used in practice, but it may lead to suboptimal precision when the weight distribution is non-uniform, as many grid points may be underutilized.
Non-uniform Quantization Grid
Non-uniform grids allocate grid points more flexibly, allowing denser spacing in high-probability regions of the weight distribution and sparser spacing in low-probability regions. This approach minimizes quantization error by adapting the grid to the data distribution. Non-uniform grids typically store the grid points in a look-up table, enabling flexible placement that better represents the original data. Figure 3 illustrates an example of affine grid and non-uniform grid applied to the weights of Llama-3-8B.
Grouped Quantization
The quantization grid for a set of weights is determined by the range $W_{\mathrm{min}},W_{\mathrm{max}}]$ within the group. Smaller group sizes allow for a narrower dynamic range, leading to finer granularity in the quantization grid and higher precision. Grouping contiguous weights into blocks is a common practice in quantization literature (awq; gptq) and ensures a balance between memory efficiency and precision.
Appendix B Controlling Range Shrinkage
In Equation 7, we enumerate candidate pairs (S,Z)(S,Z)—scaling factors and zero-points—to determine the optimal loss-error-aware affine quantization grid. This process involves iteratively refining $S$ and $Z$ by reducing the maximum value $\max(\mathbf{w})$ and increasing the minimum value $\min(\mathbf{w})$. However, excessive shrinking of the range may result in poor representation of extreme values, leading to model quality degradation.
To control the extent of range reduction, we introduce the parameter $t$, which determines the degree of shrinkage. Lower bit widths require more aggressive shrinking due to the limited number of grid points. We set $t$ for $b$-bit quantization as follows:
$t=\begin{cases}0.2T&\text{if }b=4,\\ 0.3T&\text{if }b=3,\\ 0.4T&\text{if }b=2.\end{cases}$ (8)
Appendix C LeanQuant-Exact
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
C.1 BERT Experiments with LeanQuant-Exact
@@
We compare the performance of BERT models (bert), quantized with OBQ (obq) and LeanQuantnu-Exact, on the SQuAD dataset (squad). We quantize the 12-layer BERT-base (bert) and the 3-layer BERT-3 variant from bert3 to 3 and 4 bits. OBQ and LeanQuant-Exact are calibrated using 1024 samples from the training set, and the F1 score is reported on the test set.
Table 7: Perplexity evaluations of Llama models under different quantization methods and bit widths. The results of GPTQ, AWQ, OmniQuant are from omniquant, and the results of SqueezeLLM are from sqllm. The official SqueezeLLM code does not support 2-bit quantization, and we report the available results from sqllm.
@@
Appendix D Robustness of LeanQuant Grids During Quantization
LeanQuant prevents drastic increase to the task loss by learning the quantization grid for better preservation of the precision of outlier inverse diagonals. However, since the not-yet-quantized weights will shift during the iterative quantization process and the quantization grid is fixed beforehand, one potential problem arises: the quantization grid may no longer be well-aligned with the outliers after certain iterations. Fortunately, this is not a problem in practice. The loss-error-awareness property of LeanQuant grids prevents high-norm weight perturbations $\bm{\delta}_{i}$ (Equation 3) from ocurring, hence the weights do not shift by much during the iterations. Furthermore, no new inverse-diagonal outliers will arise during the iterative quantization process. In OBQ, the inverse Hessian is updated after each iteration as follows,
where $\mathbf{H}^{-1}_{-i,-i}$ is the inverse Hessian with its $i$-th row and column removed. The remaining inverse diagonals only decrease in magnitude towards zero after each column and row removal.
Appendix E Experiment Details
Baseline Reproduction
Perplexity Evaluations
We follow the perplexity evaluation procedure described by (gptq): sequences from the test set of the WikiText2 and C4 datasets (wikitext2; c4) are concatenated into 128 sequences of length 2048 tokens for perplexity testing.
Accuracy Evaluations
We use lm-evaluation-harness (lm-eval) for evaluating zero-shot accuracy on tasks. The task names we evaluate are lambada, ai2_arc, winogrande, piqa, hellaswag, mmlu.
Appendix F Perplexity Evaluations
The perplexity evaluation results on WikiText2 (wikitext2) and C4 (c4) for quantized models are presented in Table 7.
Appendix G LLM-as-a-Judge
Appendix H More Accuracy Results
The zero-shot accuracy results on benchmarks for quantized LLaMA-7B, LLaMA-13B, Llama-2-7B (llama; llama2) are presented in Table 10. We also compare affine LeanQuant with the rotation-based quantization algorithm QuaRot (quarot), with results presented in Table 8. Furthermore, we compare affine, group-wise quantization using LeanQuantaff, OmniQuant, and AWQ in Table 9.
Table 8: Zero-shot accuracy comparison between LeanQuantaff and rotation-based quantization method QuaRot.
Method Bits ARC LAMBADA MMLU Avg.
Easy Chg Std OpenAI STEM Human. Social Other
QuaRot-RTN 4.00 65.92 46.93 65.75 71.92 49.67 51.24 65.75 64.24 60.18
Llama-3-8B QuaRot-GPTQ 4.00 76.81 50.43 67.53 74.19 51.60 53.28 70.26 68.23 64.04
LeanQuantaff 4.00 76.30 50.51 68.27 75.90 51.51 53.43 70.69 68.43 64.38
QuaRot-RTN 4.00 74.54 47.44 66.19 74.17 39.74 41.74 53.04 51.95 56.10
Llama-2-13B QuaRot-GPTQ 4.00 76.35 49.23 69.82 76.48 41.29 47.01 58.99 57.84 59.63
LeanQuantaff 4.00 76.43 49.06 69.90 76.48 42.53 47.86 60.03 58.93 60.15
Table 9: Zero-shot accuracy of affine, group-wise quantized models using LeanQuantaff, OmniQuant, and AWQ.
Method Group Size Bits ARC LAMBADA MMLU Avg.
Easy Chg Std OpenAI STEM Human. Social Other
Llama-2-7B
OmniQuant 128 4.25 75.21 43.69 66.95 72.91 35.11 37.79 46.77 46.86 53.16
AWQ 128 4.25 75.17 43.26 67.40 72.70 34.89 37.79 46.34 45.93 52.94
LeanQuantaff 128 4.25 76.26 43.00 67.75 74.44 35.65 39.02 47.94 49.79 54.23
Llama-2-13B
OmniQuant 128 4.25 78.24 47.61 69.75 75.99 42.06 47.74 59.83 58.61 59.98
AWQ 128 4.25 78.91 46.50 70.17 76.19 41.39 46.70 59.38 55.94 59.40
LeanQuantaff 128 4.25 79.00 47.18 70.75 77.95 41.96 47.86 59.54 58.97 60.40
Llama-3.1-70B-Instruct
AWQ 128 4.25 86.62 62.12 72.09 75.70 74.88 80.60 87.07 83.71 77.85
LeanQuantaff 128 4.25 86.49 61.69 72.70 76.07 75.61 80.79 87.55 83.71 78.08
@@ Appendix H More Accuracy Results
Non-uniform SqueezeLLM 4.05 74.92 40.61 66.87 71.99 29.12 29.86 29.38 34.86 56.55 78.29 69.38 52.89
LeanQuant 4.05 75.17 41.55 69.01 74.33 27.37 30.50 29.41 35.44 56.09 77.91 70.17 53.36
SqueezeLLM 3.02 71.46 37.88 61.71 71.05 23.66 26.99 26.16 30.54 54.89 77.86 68.59 50.07
LeanQuant 3.02 71.76 36.35 63.77 71.71 26.64 29.05 29.80 32.70 53.80 77.09 69.53 51.11
@@
Appendix I Quantization Cost and Overhead
The time cost of LeanQuant for different models and configurations are presented in Table 13. A comparison of GPU memory consumption for different quantization algorithms on different-sized LLMs is presented in Table 11.
Table 11: GPU memory consumption of quantization algorithms on different-sized LLMs. “OOM” means out of memory.
Llama-3-8B Llama-3-70B Mistral-Large (123B) Llama-3.1-405B
OmniQuant 25.3 GB OOM OOM OOM
SqueezeLLM OOM OOM OOM OOM
GPTQ 7.9 GB 17.1 GB 32.8 GB OOM
LeanQuant 7.9 GB 17.2 GB 33.0 GB 65.4 GB
Appendix J Inference Efficiency of Quantized Models
Table 12 presents the inference efficiency of 4-bit quantized Llama-3-8B during the decoding and prefill phases. For non-uniform LeanQuant models, we have developed a dedicated CUDA kernel for efficient inference, and we compare its efficiency against the SqueezeLLM kernel in Table 12. For affine LeanQuant models, we leverage the exllamav2 kernels (exllamav2).
The inference efficiency in Table 12 is evaluated on an NVIDIA A100-40GB GPU. For decoding, we report tokens per second per batch while generating 4096 tokens. For the prefill phase, we measure time to first token using a 4096-token prompt.
Table 12: Inference efficiency of 4-bit quantized Llama-3-8B in the decoding and prefill phases. Decoding efficiency is measured in tokens per second per batch for generating 4096 tokens, while prefill efficiency is evaluated by time to first token for a 4096-token prompt. All results are obtained on an NVIDIA A100-40GB GPU.
Decoding (tokens/s per batch) Prefill (time to first token)
Batch Size 1 Batch Size 4 Batch Size 16 Batch Size 1 Batch Size 4 Batch Size 16
SqueezeLLM 26.09 19.68 7.22 153.26s 333.36s OOM
LeanQuantnu 23.73 18.28 9.02 0.36s 1.34s 5.24s
LeanQuantaff 38.30 28.78 10.20 0.33s 1.23s 4.98s
@@ Appendix I Quantization Cost and Overhead
Appendix K Ablation Study
@@
Appendix L Loss Error Comparison
@@
Appendix M More Related Works
Quantization for Large Language Models Quantization reduces the precision of LLM parameters to achieve model compression and enable more memory-efficient inference. Calibration-free quantization approaches, such as LLM.int8 (llm_int8), NormalFloat (qlora), and Student Float (student_t), perform zero-shot quantization without requiring calibration data. In contrast, methods like GPTQ (gptq), AWQ (awq), OmniQuant (omniquant), SpQR (spqr), SqueezeLLM (sqllm), QUIP (quip), AQLM (aqlm), and QUIP# (quipsharp) leverage calibration to improve quantization quality by adapting to input data distributions. Some methods (smoothquant) extend quantization to both model weights and intermediate activations. Some approaches combine quantization-aware training to push the limits of quantization; for example, LLM-QAT (llm_qat) fine-tunes quantized models to recover model quality, BitNet (era_of_1bit) explores ternary-valued LLMs, while OneBit (onebit) demonstrates the feasibility of 1-bit quantization for LLMs.
Efficient LLMs Beyond quantization, various techniques have been proposed to enhance LLM efficiency. KV cache compression methods such as KIVI (kivi) and CQ (cq) reduce memory overhead by compressing key-value cache during LLM decoding. Pruning approaches, such as SparseGPT (sparsegpt), remove model parameters in a structured or un-structured manner to create sparse, efficient models. Model sketching (sketch_to_adapt) enables efficient fine-tuning by compressing LLMs and make them directly fine-tunable. Hardware-aware optimizations, including FlashAttention (flashattention) and NoMAD-Attention (nomad_attention), improve memory and compute efficiency for modern accelerators. Optimizer-state compression techniques (galore; i3s) reduce memory usage during pretraining and fine-tuning.
Uniform and Non-uniform Quantization Quantization techniques can be broadly categorized into uniform (affine) and non-uniform methods. Uniform quantization (uniform_quant; gptq) divides the range of values into equal-sized intervals, which is hardware-efficient but often fails to accommodate the non-uniform distribution of the weights of deep neural networks. Non-uniform quantization improves model compression by allocating precision dynamically based on data distribution. Additive Powers-of-Two Quantization (nu_powers_of_two) introduces an efficient non-uniform discretization scheme that leverages power-of-two representations. Nonuniform-to-uniform quantization (nu_to_uniform) bridges the gap between non-uniform and uniform quantization using a generalized straight-through estimator for training. NUPES (nu_nupes) formulates non-uniform post-training quantization as a power exponent search problem. Mr.BiQ (nu_mr_biq) focuses on reducing reconstruction error through post-training non-uniform quantization, improving model performance. Methods such as non-uniform step size quantization (nu_step_size) refine quantization granularity to enhance accuracy, while learning-based approaches (nu_learning_step_sizes) adaptively determine step sizes to optimize neural network quantization.